# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ 4.3. Discussion
Role of Personal Memory
Personal memory bridges short-term intent and long-term preference modeling. It fulfills three closely related roles. First, it supports preference evolution, allowing the agent to track how user preferences change across sessions rather than relying solely on fixed trait representations. Second, it provides personalized contextual grounding, preserving fine-grained user details that are essential for accurate personalization but cannot be fully encoded in static profiles. Third, it enables preference-conditioned generation, guiding the agent to interpret new queries and produce responses informed by retrieved user-relevant signals. Together, these roles transform memory from a passive log of interactions into an active component of the personalization loop.
Challenges
Despite its importance, personal memory faces several challenges. First, user preferences evolve at different rates, requiring memory systems to balance rapid adaptation with the preservation of long-term traits. Second, memory stores may accumulate redundancy or hallucinations, especially when updates rely on LLM-driven summarization or relation extraction, leading to degraded personalization over long horizons. Third, ensuring that retrieved content is not only semantically relevant but also truly aligned with user preferences remains difficult. Finally, personal memory raises privacy and transparency concerns, highlighting the need for user-controllable and privacy-preserving memory mechanisms.
@@
Memory modules also have inherent limitations. Textual memories scale poorly and require repeated rewriting; structured memories provide controllability but depend on costly extraction and risk losing contextual nuance; update mechanisms oscillate between shallow similarity rules and computationally heavy inference; and retrieval pipelines remain sensitive to ranking biases and distribution shifts. These limitations point to several broad future directions across the memory pipeline, including hybrid memory architectures that combine the semantic richness of textual representations with the organizational clarity of structured formats, more principled and user-controllable update mechanisms that ensure reliability and transparency, and improved retrieval strategies that better integrate personal alignment with semantic relevance.
5. Planning in Personalized LLM-powered Agents
Planning constitutes the cognitive engine of intelligent agents, enabling the decomposition of high-level objectives into actionable steps and the formulation of execution strategies (Luo et al., 2025a; Sun et al., 2023). In the context of personalized agents, planning transcends generic task completion; it must optimize for user utility, aligning decision paths with individual preferences, constraints, and latent intents (Zhang et al., 2024e; Gao et al., 2024a; Han et al., 2025b). Unlike standard planners that prioritize efficiency or logical correctness, personalized planners navigate a constrained search space defined by the user’s unique profile and behavioral history, often requiring the integration of long-term memory and preference models (Westhäußer et al., 2025a; Huang et al., 2025c). We categorize these methodologies into two distinct paradigms based on the integration of user signals: One-shot Planning, where personalization is internalized as a prior during reasoning, and Feedback-driven Planning, where personalization emerges dynamically through interaction (Zhang et al., 2024d; Suri et al., 2025).
@@ 5.1. One-Shot Planning
One-shot personalized planning refers to generating a complete plan in a single inference pass, where a plan is represented as a structured decomposition of subgoals and decisions conditioned on user-specific constraints and latent intent. We analyze one-shot methods along two axes: how user signals are grounded into planning representations (Preference Induction and Memory Conditioning) and how plans are synthesized via compute-intensive generation (Internal Refinement).
@@
This line of work constructs a planning-ready user state by inferring latent intent and translating heterogeneous user signals into explicit constraints or soft objectives that guide plan generation (Han et al., 2025b). For instance, ALIGNXPLORE (Li et al., 2025d) utilizes a User-description Chain-of-Thought framework to synthesize a compact, semantically grounded preference description from sparse interaction traces. Instead of retrieving raw logs, the model performs extended inductive reasoning to form global preference constraints that condition downstream planning. Similarly, in sequential recommendation, ReaRec (Tang et al., 2025) proposes a Think-Before-Recommend paradigm, using reasoning position embeddings to infer the user’s latent intent trajectory prior to plan generation. In embodied and decision-making contexts, PbP (Xu et al., 2025b) demonstrates that treating learned user preferences as intermediate abstractions substantially improves few-shot personalized planning performance.
Memory Conditioning.
Complementary to inductive constraint formation, memory-grounded methods condition one-shot planning on selectively retrieved and compressed user history, ensuring consistency with both long-term preferences and recent context (Tan et al., 2025c; Westhäußer et al., 2025a). PRIME (Zhang et al., 2025f) integrates a dual-memory architecture with a Personalized Thought Process that explicitly traces how past experiences and stable beliefs influence current decisions. Before plan generation, PRIME synthesizes a personalized thought trace via self-distillation, effectively grounding planning in the user-specific cognitive context. PersonaAgent (Zhang et al., 2025c) adopts a retrieval-augmented persona mechanism, dynamically constructing system prompts from a unified memory bank to maintain cross-session planning consistency. Related work in recommendation and dialogue systems similarly shows that structured user memory retrieval significantly improves alignment and coherence in one-shot decision-making (Huang et al., 2025c; Chen, 2025).
Internal Refinement.
Beyond direct one-pass generation, internal refinement improves it by constructing an initial plan and then internally critiquing and revising it against a user-conditioned objective. Related self-refinement templates such as Reflexion and structured critique-and-revise frameworks provide general mechanisms for generating reflective feedback and triggering plan edits (Shinn et al., 2023; Gou et al., 2023), while recent test-time-compute reasoning models make deeper within-pass self-correction increasingly practical (Muennighoff et al., 2025; Guo et al., 2025). Representative systems instantiate this paradigm in several ways. In personalized travel planning, TravelPlanner++ (Singh et al., 2024) and PTS (Shao et al., 2025) study reflection-style itinerary refinement, where intermediate plans are iteratively revised under derived user implicit long-horizon preferences. PREFINE (Ueda and Takayanagi, 2025) constructs a pseudo-user critic and user-specific rubrics from interaction history, then critiques and revises candidates to better satisfy the target user. In embodied command execution, IoTGPT (Yu et al., 2026) applies self-correction by testing intermediate command sequences in a simulated environment and revising the plan before deployment.
@@ 5.2. Feedback-Driven Planning
Feedback-driven planning treats the initial plan as a hypothesis to be refined through interaction. These systems recognize that user intent is often underspecified, evolving, or partially latent (Liu et al., 2025d; Zhang et al., 2024d), and employ iterative loops to converge on an optimal personalized solution (Suri et al., 2025; Wang et al., 2023b). We further organize feedback-driven planning by the source of feedback and how it updates the plan, including information-seeking clarification, mixed-initiative user revision, and non-user feedback signals.
@@
A central challenge in feedback-driven personalized planning lies in deciding when to ask for additional information: excessive clarification increases user burden, while premature commitment risks misinterpreting user-specific constraints or latent intent. Prior studies indicate that user requests in planning tasks are frequently underspecified or ambiguous (Liu et al., 2025d), motivating approaches that treat clarification as decision-making under uncertainty. For instance, SAGE-Agent (Suri et al., 2025) formulates selective questioning as a POMDP and issues clarification queries only when the expected value of information outweighs interaction cost. Similarly, ECLAIR (Murzaku et al., 2025) focuses on detecting missing arguments or ambiguous entities in user instructions and triggers targeted follow-up questions before committing to a plan. Additionally, Ask-before-Plan style frameworks (Zhang et al., 2024d) explicitly decouple clarification from plan synthesis, showing that resolving key uncertainties upfront improves robustness and downstream plan quality. Collectively, these methods establish clarification as a principled and integral component of feedback-driven personalized planning rather than an ad hoc interaction heuristic.
@@
User edits provide a high-fidelity personalization signal that can update user-conditioned objectives, constraints, and preferences for subsequent planning and generation. Gao et al. (2024a) show that learning latent preferences from historical edits can reduce future edit effort and improve user-specific alignment in subjective settings. Building on this mixed-initiative view, AIPOM (Kim et al., 2025a) represents plans as editable structures, enabling users to revise intermediate plans while the agent updates constraints and priorities accordingly. Similar revision loops appear in domain systems such as Dango (Chen et al., 2025a), which supports iterative correction of agent-produced data-processing workflows, and PlanFitting (Shin et al., 2025), which refines personalized exercise plans through conversational revisions. This revision-driven interaction pattern is also adopted in deployed coding agents, where users review proposed changes and iteratively provide follow-up edits or instructions to steer subsequent updates (Anthropic, 2025; Github, 2025).
@@
Beyond direct user input, plans can be refined using non-user signals that improve feasibility and preference satisfaction without repeatedly burdening the user. One line of work leverages agent-mediated feedback: VAIAGE (Liu et al., 2025a) employs specialized agents to negotiate route feasibility and recommendations under user preferences, while Collab-Rec (Banerjee et al., 2025) supports multi-stakeholder planning with explicit user-advocacy mechanisms. Another line exploits environment-mediated feedback by simulating intermediate outcomes. For instance, SceneWeaver (Yang et al., 2025) follows a Reason–Act–Reflect loop and triggers re-planning when rendered intermediate results violate user-specified attributes.
@@ 5.3. Discussion
Role of Personalized Planning.
Personalized planning serves two complementary roles in LLM-based agents. First, it operationalizes personalization by bridging upstream user intent inference and preference modeling into downstream decision-making (Chen, 2025), instantiating user-contingent objectives, constraints, and trade-offs that govern a plan. Second, particularly in feedback-driven settings (Kim et al., 2025a), planning functions as a closed-loop optimization process that alternates between eliciting informative feedback and revising the plan, enabling the agent to proactively and progressively uncover user intent and converge toward a higher degree of personalization.
Challenges.
Personalized planning raises several challenges. First, user intent and preferences are often partially observed and evolve over time (Suri et al., 2025; Liu et al., 2025d), complicating when to infer, elicit, and commit. Second, personalized utility is also multi-objective (Shao et al., 2025), requiring arbitration among competing constraints and idiosyncratic trade-offs. Moreover, interactive refinement must remain controllable and convergent, limiting user burden while avoiding unstable revisions. Existing paradigms involve inherent trade-offs. One-shot planning internalizes user signals as priors for low-latency, coherent plans, but can be brittle under sparse or shifting signals. Feedback-driven planning improves alignment via iterative elicitation and revision, at the cost of additional latency and user effort (Liu et al., 2025a).
@@
Future work may benefit from developing uncertainty-aware planning workflows that adaptively allocate internal computation and revision effort based on confidence in inferred preferences and task stakes (Muennighoff et al., 2025), combining the efficiency of one-shot planning with the alignment benefits of feedback-driven refinement. Besides, a promising direction is to learn user-conditioned planning patterns at both inter- and intra-user levels (Qiu et al., 2025a), capturing systematic differences in decomposition style and trade-off resolution while modeling their stability and drift over time, therefore providing reusable personalization signals that generalize across tasks. Finally, several aspects remain comparatively under-explored, including robustness to variations in user specification (Suri et al., 2025) and privacy-preserving planning (Wu et al., 2025e) that supports long-horizon user conditioning without exposing raw personal data.
6. Action Execution in Personalized LLM-powered Agents
Action execution is the stage where a personalized agent realizes decisions through concrete operations in the external environment and closes the loop with execution feedback (Shen, 2024; Kim et al., 2024). Unlike purely task-centric execution, this layer must ground actions in user-specific constraints and preferences, and remain adaptive when tools or environments produce unexpected outcomes. We organize personalized action execution into two stages: a Pre-action Stage that selects and parameterizes actions under user-conditioned constraints, and a Post-action Stage that leverages execution signals for recovery and preference-consistent outcome realization.
@@ 6.1. Pre-action Stage
The pre-action stage focuses on user-conditioned action decision and grounding. It translates a selected intent or plan step into executable tool calls or environment operations under user-specific constraints. We further decompose this stage into action policy, which selects among functionally valid execution options for the same step, and action grounding, which instantiates the selected action with user-specific arguments and realizations.
@@ 6.1.1. Action Policy
At this level, personalization manifests as a bias over execution choices, spanning both tool utilization and higher-level action decisions. PEToolLLaMA (Xu et al., 2025c) is an early effort that formalizes personalized tool learning from interaction history, constructs a tool-use dataset with diverse user preferences, and trains models to improve personalized tool selection. In the web-agent setting, PUMA (Cai et al., 2025) adopts a similar preference-aware learning strategy, showing that personalization can also shape action choices during web navigation and task completion. ToolSpectrum (Cheng et al., 2025) further incorporates user profiles together with environmental context, indicating that context-aware conditioning can improve user-aligned tool utilization. Beyond selecting appropriate tools, ColorAgent (Li et al., 2025f) learns to proactively engage the user to clarify ambiguous intentions or incomplete instructions. Moving beyond tool invocation, ValuePilot (Luo et al., 2025b) studies action-level decision making under user values, where agents evaluate and select actions according to user value preferences to produce context-sensitive and individualized choices.
@@ 6.1.2. Action Grounding
Action grounding operationalizes a selected action by instantiating it into executable tool calls with user-conditioned constraints and parameters (Cai et al., 2025; Cheng et al., 2025). While many requests can be grounded by directly applying user-specific settings, personalization becomes more challenging in corner cases. A frequent grounding-time failure mode is infeasibility, where the instantiated query becomes unsatisfiable under the imposed constraints. AWARE-US (Kurmaz, 2026) frames this as a preference-aware query repair problem, arguing that agents should restore feasibility by relaxing the least preferred constraints to the user rather than applying default heuristics. Another recurring issue is the omission of essential tool arguments. Huang et al. (2025d) study how agents infer unspecified arguments from user profiles, reducing execution friction while maintaining alignment with user intent.

### v2

@@ 4.3. Discussion
Role and Challenges
Personal memory bridges short-term interaction context with long-term preference modeling, fulfilling three closely related roles. It provides personalized contextual grounding by preserving fine-grained user details that cannot be fully encoded in static profiles, supports preference evolution by allowing the agent to track how user preferences change across sessions, and enables preference-conditioned generation by supplying user-relevant signals at inference time. At the same time, these roles make personal memory difficult to design effectively. User preferences evolve at different rates, requiring memory systems to balance rapid adaptation with the preservation of long-term traits. Memory stores may also accumulate redundancy or hallucinations, especially when updates rely on LLM-based summarization or relation extraction, leading to degraded personalization over long horizons. Moreover, retrieving content that is not only semantically relevant but also truly aligned with user preferences remains a persistent challenge. Finally, personal memory raises privacy and transparency concerns, highlighting the need for user-controllable and privacy-preserving memory mechanisms.
@@
These limitations point to several promising directions across the memory pipeline. One is hybrid memory architectures that combine the semantic richness of textual representations with the organizational clarity of structured formats. Another is more principled and user-controllable update mechanisms that improve reliability, transparency, and long-term consistency. A third is retrieval strategies that better integrate personal alignment with semantic relevance, so that retrieved information is not only related to the query but also truly useful for personalization.
5. Planning
In PLAs, planning is the stage at which user-specific information is translated into actionable decisions. While profile modeling captures what is known about the user and memory preserves user-related information over time, planning determines how these signals shape subgoal decomposition, strategy selection, and trade-off resolution during task execution (Luo et al., 2025a; Sun et al., 2023). Thus, it goes beyond generic task completion and instead seeks to optimize user utility by aligning decisions with individual preferences, constraints, and latent intents (Zhang et al., 2024e; Gao et al., 2024a; Han et al., 2025a). We organize existing methods into two broad paradigms according to how user signals enter the planning process: One-shot Planning, in which personalization is incorporated as a prior during plan generation, and Feedback-driven Planning, in which personalization is progressively refined through interaction (Zhang et al., 2024d; Suri et al., 2025).
@@ 5.1. One-Shot Planning
One-shot personalized planning generates a complete plan within a single inference pass, where the plan is represented as a structured decomposition of subgoals and decisions conditioned on user-specific constraints and latent intent. We analyze this paradigm along two complementary axes: user signal grounding, which concerns how user signals are incorporated into the planning process, and internal refinement, which concerns how an initial plan is internally improved within the same inference pass.
5.1.1. User Signal Grounding
One-shot planning can ground user-specific information in two main ways. One directly conditions planning on existing user representations, such as explicit profiles, retrieved memory, or compressed interaction history. The other infers planning-ready constraints or objectives from heterogeneous user signals before plan generation.
Profile and Memory Conditioning
These methods condition one-shot planning on existing user representations, such as explicit profiles, retrieved memory, or compressed interaction history, thereby helping maintain consistency with both long-term preferences and recent context (Tan et al., 2025c; Westhäußer et al., 2025a). PRIME (Zhang et al., 2025d) integrates a dual-memory architecture with a Personalized Thought Process that explicitly traces how past experiences and stable beliefs influence current decisions. Before plan generation, PRIME synthesizes a personalized thought trace via self-distillation, effectively grounding planning in the user-specific cognitive context. PersonaAgent (Zhang et al., 2025c) adopts a retrieval-augmented persona mechanism, dynamically constructing system prompts from a unified memory bank to maintain cross-session planning consistency. Related work in recommendation and dialogue systems similarly shows that structured user memory retrieval significantly improves alignment and coherence in one-shot decision-making (Huang et al., 2025b; Chen, 2025).
@@ 5.1.1. User Signal Grounding
Complementary to direct conditioning, this line of work constructs a planning-ready user state by inferring latent intent and translating heterogeneous user signals into explicit constraints or soft objectives that guide plan generation (Han et al., 2025a). For instance, ALIGNXPLORE (Li et al., 2025e) utilizes a User-description Chain-of-Thought framework to synthesize a compact, semantically grounded preference description from sparse interaction traces. Instead of retrieving raw logs, the model performs extended inductive reasoning to form global preference constraints that condition downstream planning. Similarly, in sequential recommendation, ReaRec (Tang et al., 2025) proposes a Think-Before-Recommend paradigm, using reasoning position embeddings to infer the user’s latent intent trajectory prior to plan generation. In embodied and decision-making contexts, PbP (Xu et al., 2025b) demonstrates that treating learned user preferences as intermediate abstractions substantially improves few-shot personalized planning performance.
5.1.2. Internal Refinement
Internal refinement improves one-shot planning by first generating an initial plan and then internally critiquing and revising it against a user-conditioned objective. General self-refinement frameworks, such as Reflexion and structured critique-and-revise methods, provide reusable mechanisms for reflective feedback and plan editing (Shinn et al., 2023; Gou et al., 2023), while recent test-time-compute reasoning models make deeper within-pass self-correction increasingly practical (Muennighoff et al., 2025; Guo et al., 2025). Representative systems instantiate this paradigm in several domains. In personalized travel planning, TravelPlanner++ (Singh et al., 2024) and PTS (Shao et al., 2025) iteratively refine intermediate itineraries under implicit long-horizon user preferences. PREFINE (Ueda and Takayanagi, 2025) constructs a pseudo-user critic and user-specific rubrics from interaction history to critique and revise candidate plans. In embodied command execution, IoTGPT (Yu et al., 2026) performs self-correction by testing intermediate command sequences in a simulated environment before deployment.
@@ 5.2. Feedback-Driven Planning
Feedback-driven planning treats an initial plan as a provisional hypothesis to be refined through interaction. It is motivated by the fact that user intent is often underspecified, evolving, or only partially observable (Liu et al., 2025d; Zhang et al., 2024d). Rather than committing to a fixed plan in a single pass, these systems iteratively update plans to converge toward a more personalized solution (Suri et al., 2025; Wang et al., 2023b). We organize this paradigm according to the source of feedback and how it updates the plan, including clarification seeking, user revision, and non-user feedback.
@@
A central challenge in feedback-driven personalized planning lies in deciding when to ask for additional information: excessive clarification increases user burden, while premature commitment risks misinterpreting user-specific constraints or latent intent. Prior studies indicate that user requests in planning tasks are frequently underspecified or ambiguous (Liu et al., 2025d), motivating approaches that treat clarification as decision-making under uncertainty. For instance, SAGE-Agent (Suri et al., 2025) formulates selective questioning as a POMDP and issues clarification queries only when the expected value of information outweighs interaction cost. Similarly, ECLAIR (Murzaku et al., 2025) focuses on detecting missing arguments or ambiguous entities in user instructions and triggers targeted follow-up questions before committing to a plan. Additionally, Ask-before-Plan style frameworks (Zhang et al., 2024d) explicitly decouple clarification from plan synthesis, showing that resolving key uncertainties upfront improves robustness and downstream plan quality.
@@
User edits provide a high-fidelity personalization signal that can update user-conditioned objectives, constraints, and preferences for subsequent planning and generation. Gao et al. (2024a) shows that learning latent preferences from historical edits can reduce future edit effort and improve user-specific alignment in subjective settings. Building on this mixed-initiative view, AIPOM (Kim et al., 2025a) represents plans as editable structures, enabling users to revise intermediate plans while the agent updates constraints and priorities accordingly. Similar revision loops appear in domain systems such as Dango (Chen et al., 2025a), which supports iterative correction of agent-produced data-processing workflows, and PlanFitting (Shin et al., 2025), which refines personalized exercise plans through conversational revisions. Similar revision patterns are also increasingly visible in deployed coding assistants (Anthropic, 2025; Github, 2025).
@@
Beyond direct user input, plans can also be refined through non-user signals that improve feasibility and preference satisfaction without repeatedly increasing interaction burden. One source is agent-mediated feedback. For example, VAIAGE (Liu et al., 2025a) employs specialized agents to negotiate route feasibility and recommendations under user preferences, while Collab-Rec (Banerjee et al., 2025) supports multi-stakeholder planning through explicit user-advocacy mechanisms. Another source is environment-mediated feedback, where intermediate outcomes are simulated and used to trigger re-planning. For instance, SceneWeaver (Yang et al., 2025) follows a Reason–Act–Reflect loop and revises plans when intermediate rendering results reveal violations.
@@ 5.3. Discussion
Role and Challenges.
Planning serves two complementary roles in PLAs. First, it operationalizes personalization by translating user intent inference and preference modeling into downstream decisions, instantiating user-contingent objectives, constraints, and trade-offs that govern plan generation (Chen, 2025). Second, especially in feedback-driven settings (Kim et al., 2025a), planning functions as a closed-loop optimization process that alternates between eliciting informative feedback and revising the plan, enabling the agent to progressively uncover user intent and move toward stronger personalization. However, these roles introduce several challenges at the same time. User intent and preferences are often only partially observed and may evolve over time (Suri et al., 2025; Liu et al., 2025d), making it difficult to decide when to infer, elicit, and commit. Moreover, personalized utility is inherently multi-objective (Shao et al., 2025), requiring arbitration among competing constraints and idiosyncratic trade-offs. In interactive settings, refinement must also remain controllable and convergent, limiting user burden while avoiding unstable revisions. These difficulties are reflected in the trade-off between existing paradigms: one-shot planning offers low-latency and coherent decisions by internalizing user signals as priors, but can be brittle under sparse or shifting signals, whereas feedback-driven planning improves alignment through iterative elicitation and revision at the cost of additional latency and user effort (Liu et al., 2025a).
@@
Promising directions include uncertainty-aware planning workflows that adaptively allocate internal computation and revision effort according to confidence in inferred preferences and task stakes (Muennighoff et al., 2025), thereby combining the efficiency of one-shot planning with the alignment benefits of feedback-driven refinement. Another direction is to learn user-conditioned planning patterns at both inter- and intra-user levels (Qiu et al., 2025a), capturing systematic differences in decomposition style and trade-off resolution while modeling their stability and drift over time. More broadly, important but still under-explored directions include robustness to variation in user specification (Suri et al., 2025) and privacy-preserving planning (Wu et al., 2025f) that supports long-horizon user conditioning without exposing raw personal data.
6. Action Execution
Action execution is the stage at which a PLA realizes decisions through concrete operations in external environments and receives feedback from execution outcomes (Shen, 2024; Kim et al., 2024). Unlike purely task-centric execution, personalized execution must ground actions in user-specific constraints and preferences, while remaining adaptive when tools or environments produce unexpected outcomes. We organize personalized action execution into two stages: a pre-action stage that selects and parameterizes actions under user-conditioned constraints, and a post-action stage that leverages execution signals for recovery and preference-consistent outcome realization.
@@ 6.1. Pre-action Stage
The pre-action stage focuses on user-conditioned action decision and grounding. It translates a selected intent or plan step into executable tool calls or environment operations under user-specific constraints. We further decompose this stage into action policy, which selects among functionally valid execution options for a given step, and action grounding, which instantiates the selected action with user-specific arguments and realizations.
@@ 6.1.1. Action Policy
At this level, personalization manifests as a bias over functionally valid execution choices, spanning both tool utilization and higher-level action selection. PEToolLLaMA (Xu et al., 2025c) formalizes personalized tool learning from interaction history and trains models to improve preference-aware tool selection. In web-agent settings, PUMA (Cai et al., 2025) adopts a similar preference-aware learning strategy, while ToolSpectrum (Cheng et al., 2025) further conditions tool use on both user profiles and environmental context. Beyond tool invocation, ValuePilot (Luo et al., 2025b) studies action-level decision making under user values, showing that execution choices can be guided by individualized value preferences. In cases of ambiguity, action policy may also involve proactive clarification. For example, ColorAgent (Li et al., 2025g) learns to engage the user when intentions or instructions are incomplete before committing to an execution choice.
@@ 6.1.2. Action Grounding
Action grounding operationalizes a selected action by instantiating it into executable tool calls with user-conditioned constraints and parameters (Cai et al., 2025; Cheng et al., 2025). While many requests can be grounded by directly applying user-specific settings, personalization becomes more challenging in corner cases. A frequent grounding-time failure mode is infeasibility, where the instantiated query becomes unsatisfiable under the imposed constraints. AWARE-US (Kurmaz, 2026) frames this as a preference-aware query repair problem, arguing that agents should restore feasibility by relaxing the least preferred constraints to the user rather than applying default heuristics. Another recurring issue is the omission of essential tool arguments. Huang et al. (2025c) study how agents infer unspecified arguments from user profiles, reducing execution friction while maintaining alignment with user intent.